OpenStudy (anonymous):

What is the midpoint of (3, 4) and (–1, –2)? (1, 1) (2, 1) (0, 0) (0, 3)

OpenStudy (aroub):

midpoint formula: (x1+x2/2, y1+y2/2)

OpenStudy (aroub):

=> (3-1/2, 4-2/2) = (2/2,2/2) = (1,1)

OpenStudy (anonymous):

so is b

OpenStudy (aroub):

Why?

OpenStudy (anonymous):

so is a

OpenStudy (aroub):

yeah :)

OpenStudy (anonymous):

Thank U

OpenStudy (anonymous):

read your textbook to find out why the mid-point formula is the way it is.

OpenStudy (anonymous):

i did

OpenStudy (aroub):

You're welcome!